the school has a ractangular playground. the ratio of the length of the playground to its width is 3:4. the length of the playground is 45 yards. find the perimeter and area.
Answer by CubeyThePenguin(3113)

length : width = 3 : 4 = 45 : x

width = x = (4 * 45)/3 = 60

perimeter = 2(60) + 2(45) = 210 yd

area = (60)(45) = 2700 yd^2